FIXDLSRPS(1) FIXDLSRPS(1) NAME fixdlsrps−filtertofixDviLaser/PSdocumentstoworkwithPSUtils SYNOPSIS fixdlsrps<dvilaser.ps>Fixed.ps DESCRIPTION Fixdlsrpsisaperlfilterwhich"fixes"PostScriptgeneratedfromtheDviLaser/PSdriversothatit workscorrectlywithAngusDuggan’spsutilspackage. AUTHOR Copyright(C)AngusJ.C.Duggan1991-1995 SEEALSO psbook(1),psselect(1),pstops(1),epsffit(1),psnup(1),psresize(1),fixfmps(1),fixmacps(1),fixpsditps(1),fixpspps(1),fixtpps(1),fixwfwps(1),fixwpps(1),fixscribeps(1),fixwwps(1),fixdlsrps(1), extractres(1),includeres(1),psmerge(1) TRADEMARKS PostScriptisatrademarkofAdobeSystemsIncorporated. PSUtilsRelease1Patchlevel17 1